y/b=3/2 and b/r=3/5 how many r's would there be in 255? please email me the answer and how to solve it.
Answer by richwmiller(17219)

y/b=3/2,
b/r=3/5,
255/r=x
x = 153/b,
There are 153/b r's in 255
r = (5 b)/3, b!=0, y = (3 b)/2